40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6338     Entry time: Wed Feb 29 01:02:06 2012
Author: Den 
Type: Update 
Category: PEM 
Subject: seis box measured 

I've measured the input signal to the seismic box from seismometer Guralp 1. The spectrum of the signal in the "input +" (TP 1) is

input.png

 

The spectrum below 1 Hz is ~250 uV/sqrt(Hz). As the input is differential, then the input voltage is 0.5 mV/sqrt(Hz). The spectrum of the "output +" signal (TP 2) is

output.png

 

So the gain at ~ 1Hz is ~20. I've measured the transfer function between the "input +" and "output +" (TP1 and TP2) for all 9 circuits

tf.png

The channels 1-6 are of new modification and have gain ~20 at the frequencies 0.2 - 100 Hz. Below 0.2 Hz the gain is reduced. 100 Hz - cut off frequency of the low-pass filters. Meanwhile channels 7-9 (old configuration) have much more gain and 10_50 Hz filters work here.

The coherence between  "input +" and "output +" (TP1 and TP2) for 9 circuits is

 

coherence.png

We can see that channel VERT 3 is very bad. For others coherence is lost below 0.2 Hz. The spectrum analyzer noise measured is ~1000 times less then the signal at these frequencies. I'll pay more attention to this loss of coherence at low frequencies. Something is noisy.

ELOG V3.1.3-